# Rethinking industrial artificial intelligence: a unified foundation framework

### v1

@@ 4.3.1 Algorithm design
Algorithm design serves as the foundation for developing AI models. Developers should select or design models that align with data characteristics, domain constraints, and application requirements. Typical options range from traditional machine learning and statistical methods—such as decision tree, random forest, support vector machine, gradient boosting, k-nearest neighbors, Gaussian mixture model, and various clustering algorithms—to deep learning architectures, including convolutional neural networks (CNNs), recurrent neural networks (RNNs), graph neural networks (GNNs), Transformers, and their variants [5, 46]. These algorithms are applicable to problems involving both continuous variables (e.g., vibration signals, temperature, pressure) and discrete variables (e.g., operational states, fault modes, control events). Beyond traditional machine learning and deep learning architectures, physics-informed neural networks (PINNs) have emerged as valuable tools in industrial AI [62, 33]. PINNs incorporate physical laws and equations into the learning process, improving model reliability in scenarios where data is sparse but prior knowledge is available [62, 33]. Additionally, multi-modal learning approaches, which combine information from multiple modalities — such as images, text, audio, tabular data, and sensor measurements — enable the development of more comprehensive models suitable for complex industrial scenarios [13, 53]. Furthermore, foundation models, which are pre-trained on large-scale diverse datasets, offer a promising approach for transferable and adaptable AI across different industrial domains. These models can be fine-tuned for specific tasks, allowing for efficient deployment with reduced training time and improved generalization [67]. In the process of algorithm design, developers may utilize prior knowledge and experience to create new model architectures or iteratively refine existing models through systematic experimentation and testing. When open-source implementations are available, they can be reproduced and further extended based on established methods.
@@ 4.3.2 Model interpretability
Model interpretability is crucial for ensuring AI models are transparent and explainable. One important direction is interpretable machine learning, which focuses on developing models that transparent by design, such as regression models, decision trees, and generalized additive models. These models, along with techniques like rule-based learning and sparse linear models, can provide clear and consistent explanations, making them suitable for industrial applications [39, 36]. However, with the increasing use of deep learning models, which are often treated as black boxes, it has become challenging to understand and explain how such models make predictions. To address this, developers are encouraged to use post-hoc explanation methods such as SHapley Additive exPlanations (SHAP) values [41], Local Interpretable Model-agnostic Explanations (LIME) [45], and integrated gradients. These methods enable feature importance analysis and provide explanations of model predictions in both global and local contexts. In addition, causal inference techniques can enhance model interpretability by identifying cause-and-effect relationships between input variables and model outcomes. Approaches such as structural causal models, counterfactual analysis, and invariant causal prediction can be used to provide more robust and stable explanations, particularly in dynamic industrial settings [38].
@@ 4.3.3 Model adaptation and generalization
Model adaptation refers to the ability of an AI model to adjust its parameters or structure to accommodate new data distributions, domains, or operational scenarios without requiring complete retraining. Meanwhile, model generalization refers to the model’s ability to maintain accurate predictions when exposed to previously unseen data or tasks. They are essential to ensure that models remain robust, reliable, and effective in real-world industrial settings. Several key techniques support model adaptation and generalization. Transfer learning and domain adaptation allow models trained in one context or domain to be fine-tuned or adjusted for use in another, reducing the need for large amounts of new labeled data [47, 35]. Federated learning enables decentralized model training across multiple industrial sites, allowing collaborative improvement of models while preserving data privacy and security [4]. Few-shot and zero-shot learning techniques allow models to make accurate predictions in new domains with very limited (few-shot) or no (zero-shot) task-specific training data [49, 61]. These are especially valuable when collecting and labeling new industrial data is time-consuming or expensive. Online learning allows models to continuously update as new data streams become available, maintaining relevance and accuracy in real-time applications [23]. Continual learning techniques enable models to incrementally incorporate knowledge from new data without catastrophic forgetting of previously learned information, supporting long-term model evolution in dynamic environments [60].
@@ 4.3.4 Model optimization and deployment readiness
Model optimization is essential for improving the performance, efficiency, and reliability of AI models in industrial applications. It involves tuning model parameters, selecting appropriate architectures, and improving computational efficiency. Common techniques include hyperparameter optimization methods such as grid search, random search, and Bayesian optimization, which systematically explore the parameter space to find optimal configurations [64]. In addition to hyperparameter tuning, model optimization also addresses structural design and computational efficiency to meet industrial requirements. Neural architecture search (NAS) methods, including differentiable NAS and evolutionary algorithms, can automate the discovery of architectures that achieve optimal trade-offs between accuracy, latency, and memory consumption [14]. Model compression techniques such as pruning (magnitude-based or structured pruning), weight quantization, and low-rank factorization are commonly used to reduce inference time and deployment costs without significant loss of performance [12]. Knowledge distillation further enhances deployment readiness by transferring knowledge from large teacher models to smaller student models [20]. Furthermore, multi-condition testing could be performed to ensure that models perform reliably across different operating scenarios [26].
@@ 4.4 Industrial AI methodology platform
The industrial AI methodology platform provides structured approaches, architectures, and guiding principles to ensure AI applications are systematically integrated into industrial systems. This platform includes multiple well-established methodologies that shape AI integration in industrial settings. Examples include: (i) Industrial large knowledge model (ILKM): ILKM bridges LLMs with domain-specific industrial knowledge to support reasoning, explanation, and contextual decision-making in industrial AI [31]. (ii) 5C cyber-physical system: CPS provides a structured approach for integrating AI with industrial systems through five levels: connection, conversion, cyber, cognition, and configuration, enabled by operation technology, analytic technology, data technology, and platform technology [29, 28]. (iii) Stream-of-quality & stream-of-X: This shows a paradigm that defines a structured methodology for continuous monitoring, optimization, and decision-making in industrial systems, particularly for multi-stage manufacturing processes [30]. (iv) Digital twin: It integrates data, simulation, and services to create a virtual representation of physical entities, achieving industrial intelligence by utilizing AI, IoT, and ML techniques [16].
@@
While existing methodologies have played an important role in advancing industrial AI, their implementation can be further enhanced by the proposed industrial AI foundation framework. Rather than redefining the core methodologies, knowledge, data, and model modules offer comprehensive and systematic views that support both existing and emerging industrial AI approaches. By promoting more structured development processes, they ensure that the design, implementation, and deployment of Industrial AI solutions are more effective, reliable, and scalable. A concrete example is using ILKM to enable intelligent question-answering (QA) systems for maintenance personnel, who frequently need to query complex maintenance manuals and troubleshooting guides during equipment servicing. The knowledge module enables structured extraction of key procedures, fault descriptions, and parameter settings from maintenance manuals and technical documentation. This content is organized into knowledge graphs and indexed repositories that can be efficiently queried by LLMs. The data module ensures that domain-specific datasets — including historical maintenance logs and annotated QA pairs — are properly documented, versioned, and linked to the knowledge base. Standardized metadata (e.g., machine type, operating conditions, fault categories) allows the LLM to retrieve relevant answers aligned with specific equipment and scenarios. The model module guides the fine-tuning of LLMs on domain-specific QA datasets. It also defines best practices for model evaluation and model optimization strategies for efficient inference. As a result, maintenance engineers can ask context-aware questions (e.g., ”What should I check if vibration levels exceed the threshold in a [specific] machine?”), and the ILKM-powered QA system provides answers grounded in both structured knowledge and validated historical data.
@@ 5 Case study
In this section, we demonstrate the application of the proposed industrial AI foundation framework through a case study on intelligent fault diagnosis for rotating machinery. We utilize a publicly available gearbox vibration dataset provided by the PHM Society [51]. The dataset consists of 2016 labeled samples collected under 78 different operating conditions, covering 7 distinct fault classes, including various gear damage levels (class 1, 2, 3, 4, 6, 8) and healthy state (class 0). Each sample comprises time-domain vibration signals recorded between 3-12 seconds at a sampling frequency of 20480 Hz. The problem addressed in this study is to evaluate the multi-class classification performance of the model using this dataset under various operating conditions. Figure 3 illustrates the workflow of applying the industrial AI foundation framework, guiding knowledge extraction, data preparation, model development, and evaluation in a structured and systematic manner.
Figure 3: Case study of applying the industrial AI foundation framework for AI-driven rotating machinery fault diagnosis.
@@
The process begins with the knowledge module. Under its guidance, domain knowledge was extracted from existing publications related to the dataset by leveraging the latest LLMs, GPT-4o, via OpenAI’s API. Researchers interact with the LLM using targeted research questions such as: “Can you provide the background information about this dataset?”, “In [target] paper, how did the authors perform data processing?”, and “How can one develop specific machine learning models such as 1D-CNN, LSTM, or Transformer for this task?” GPT-4o provides structured responses, summaries, and code generation examples, helping researchers quickly understand the dataset’s background, signal characteristics, data preparation requirements, feature extraction strategies, and model development practices.
@@
The data module is applied to ensure systematic preprocessing. A structured pipeline is designed under the guidance of domain knowledge from prior studies. The process begins with data cleaning to improve data quality, followed by data segmentation to increase the number of usable samples for model training. Next, feature engineering is performed using Fast Fourier Transform (FFT) to convert raw time-series signals into frequency domain features. The dataset is then split into training, validation, and test sets in a 3:1:1 ratio to ensure robust model evaluation. Throughout this stage, researchers continue to interact with LLMs to validate the soundness of their preprocessing strategies, or quickly obtain code examples to implement new ideas.
@@
After ensuring that the data is AI-ready, in the model module, researchers develop and evaluate eight different AI models based on previous methodologies and their own expertise. These include: (i) tree-based model (decision tree and random forest), (ii) CNN-based model (navie 1D-CNN, residual 1D-CNN), (iii) LSTM-based model (navie LSTM, bi-LSTM, hybrid-LSTM) and (iv) transformer-based model (vanilla transformer). Deep learning models were selected for their ability to automatically extract complex patterns from large-scale, frequency domain features and for their proven robustness in handling variations in operating conditions without the need for extensive manual feature engineering. Hyperparameter tuning and performance evaluation were performed following experience indicated in the previous research and researchers’ own development experience.
@@
The classification accuracy and confusion matrix are presented in Figure 3, demonstrating high performance across various architectures, with the transformer model achieving the highest accuracy of 99.47%. These results highlight the potential of advanced deep learning architectures in industrial fault diagnosis. Compared with previous studies on similar gearbox fault classification tasks, the proposed approach shows clear improvement. For instance, Su and Lee developed a residual CNN that achieved 96.99% accuracy [51], Vaerenberg et al. used power spectral density (PSD) preprocessing, log normalization, and a 3-layer CNN to reach 96.9% accuracy [57], and Gauriat et al. proposed multi-class neural additive models with 92.03% accuracy [18]. The higher performance achieved in this study is attributed to not only the model design but also the systematic application of the industrial AI foundation framework. The structured guidance from the knowledge, data, and model modules ensured consistent data preprocessing, appropriate model selection, and effective hyperparameter tuning, ultimately enhancing the reliability, scalability, and real-world applicability.
@@ 6 Future direction
To further strengthen the Industrial AI foundation framework, several key directions require further exploration. One critical aspect is talent development. Incorporating 4P-based learning (Principle, Practice, Problem-solving, Professional) and interdisciplinary training in AI/ML, engineering, and industrial applications could benefit the next generation of industrial AI practitioners. Another promising direction is data foundry, which aims to establish a standardized framework for industrial dataset collection, annotation, benchmarking, and management. A well-structured data foundry would enhance collaborative AI research, reproducibility, and cross-industry data sharing, while also facilitating the hosting of Industrial AI Data Challenge Competitions to promote innovation and benchmark AI model performance on industrial datasets. Meanwhile, the foundation framework can be extended toward discrete event dynamic systems and hybrid control systems. This would involve adapting the knowledge–data–model modules to better handle event-based transitions, symbolic representations, and hierarchical system logic. Furthermore, future research should explore the development of an LLM-assisted intelligent knowledge management system to better make use of historical case studies, domain expertise, and best practices. Such a system could autonomously acquire, structure, and retrieve relevant information, providing researchers and engineers with contextualized, actionable insights to improve AI model development and decision-making in industrial applications.
@@ 7 Conclusion
This paper presents a unified industrial AI foundation framework, structured into three core modules—the knowledge module, data module, and model module—which collectively support and enhance the industrial AI methodology platform. The role of industrial AI is explored in detail, and a case study on intelligent rotating machinery diagnosis demonstrates the framework’s potential. Furthermore, several future directions are also discussed. Overall, our unified Industrial AI foundation framework provides a systematic approach to developing, validating, and deploying industrial AI solutions for future industries. Despite these contributions, we acknowledge that real industrial data acquisition and labeling remain challenging. Furthermore, while the framework leverages LLMs for knowledge extraction, human guidance is still required, and computational costs may limit deployment in certain industrial settings. Moving forward, we encourage further research and collaboration between academia, industry, and AI practitioners to refine and expand this framework, fostering the next generation of intelligent, knowledge-data-driven industrial AI ecosystems.

### v2

@@ 4.3.1 Algorithm design
Algorithm design serves as the foundation for developing AI models. Developers should select or design models that align with data characteristics, domain constraints, and application requirements. Typical options range from traditional ML and statistical methods—such as decision tree, random forest, support vector machine, gradient boosting, k-nearest neighbors, Gaussian mixture model, and various clustering algorithms—to deep learning architectures, including convolutional neural networks (CNNs), recurrent neural networks, graph neural networks, transformers, and their variants [5, 46]. These algorithms are applicable to problems involving both continuous variables (e.g., vibration signals, temperature, pressure) and discrete variables (e.g., operational states, fault modes, control events). Beyond traditional ML and deep learning architectures, physics-informed neural networks (PINNs) have emerged as valuable tools in industrial AI [62, 33]. PINNs incorporate physical laws and equations into the learning process, improving model reliability in scenarios where data is sparse but prior knowledge is available. Additionally, multi-modal learning approaches, which combine information from multiple modalities — such as images, text, audio, tabular data, and sensor measurements — enable the development of more comprehensive models suitable for complex industrial scenarios [13, 53]. Furthermore, foundation models, which are pre-trained on large-scale diverse datasets, offer a promising approach for transferable and adaptable AI across different industrial domains. These models can be fine-tuned for specific tasks, allowing for efficient deployment with reduced training time and improved generalization [67]. In the process of algorithm design, developers may utilize prior knowledge and experience to create new model architectures or iteratively refine existing models through systematic experimentation and testing. When open-source implementations are available, they can be reproduced and further extended based on established methods.
@@ 4.3.2 Model interpretability
Model interpretability is crucial for ensuring that AI models are transparent and explainable. One important direction is interpretable ML, which focuses on developing models that transparent by design, such as regression models, decision trees, and generalized additive models. These models, along with techniques such as rule-based learning and sparse linear models, can provide clear and consistent explanations, making them suitable for industrial applications [39, 36]. However, with the increasing use of deep learning models, which are often treated as black boxes, it has become challenging to understand and explain how such models make predictions. To address this, developers are encouraged to use post-hoc explanation methods such as SHAP values [41], Local Interpretable Model-agnostic Explanations [45], and integrated gradients. These methods enable feature importance analysis and provide explanations of model predictions in both global and local contexts. In addition, causal inference techniques can enhance model interpretability by identifying cause-and-effect relationships between input variables and model outcomes. Approaches such as structural causal models, counterfactual analysis, and invariant causal prediction can be used to provide more robust and stable explanations, particularly in dynamic industrial settings [38].
@@ 4.3.3 Model adaptation and generalization
Model adaptation refers to the ability of an AI model to adjust its parameters or structure to accommodate new data distributions, domains, or operational scenarios without requiring complete retraining. Meanwhile, model generalization refers to the model’s ability to maintain accurate predictions when exposed to previously unseen data or tasks. They are essential to ensure that models remain robust, reliable, and effective in real-world industrial settings. Several key techniques support model adaptation and generalization. Transfer learning and domain adaptation allow models trained in one context or domain to be fine-tuned or adjusted for use in another, reducing the need for large amounts of new labeled data [47, 35]. Federated learning enables decentralized model training across multiple industrial sites, allowing collaborative improvement of models while preserving data privacy and security [4]. In addition, few-shot and zero-shot learning techniques allow models to make accurate predictions in new domains with limited (few-shot) or no (zero-shot) task-specific training data [49, 61]. These are especially valuable when collecting and labeling new industrial data is time-consuming or expensive. Online learning allows models to continuously update as new data streams become available, maintaining relevance and accuracy in real-time applications [22]. Likewise, continual learning techniques enable models to incrementally incorporate knowledge from new data without catastrophic forgetting of previously learned information, supporting long-term model evolution in dynamic environments [60].
@@ 4.3.4 Model optimization and deployment readiness
Model optimization is essential for improving the performance, efficiency, and reliability of AI models in industrial applications. It involves tuning model parameters, selecting appropriate architectures, and improving computational efficiency. Common techniques include hyperparameter optimization methods, such as grid search, random search, and Bayesian optimization, which systematically explore the parameter space to find optimal configurations [64]. In addition to hyperparameter tuning, model optimization also addresses structural design and computational efficiency to meet industrial requirements. Neural architecture search (NAS) methods, including differentiable NAS and evolutionary algorithms, can automate the discovery of architectures that achieve optimal trade-offs between accuracy, latency, and memory consumption [14]. Model compression techniques such as pruning (magnitude-based or structured pruning), weight quantization, and low-rank factorization are commonly used to reduce inference time and deployment costs without significant loss of performance [12]. Knowledge distillation further enhances deployment readiness by transferring knowledge from large teacher models to smaller student models [19]. Furthermore, multi-condition testing could be performed to ensure that models perform reliably across different operating scenarios [25].
@@ 4.4 Industrial AI methodology platform
The industrial AI methodology platform provides structured approaches, architectures, and guiding principles to ensure AI applications are systematically integrated into industrial systems. This platform includes multiple well-established methodologies that shape AI integration in industrial settings. Examples include: (i) ILKM: ILKM bridges LLMs with domain-specific industrial knowledge to support reasoning, explanation, and contextual decision-making in industrial AI [31]; (ii) 5C CPS: CPS provides a structured approach for integrating AI with industrial systems through five levels: connection, conversion, cyber, cognition, and configuration, enabled by operation technology, analytic technology, data technology, and platform technology [29, 28]; (iii) Stream-of-quality & stream-of-X: This shows a paradigm that defines a structured methodology for continuous monitoring, optimization, and decision-making in industrial systems, particularly for multi-stage manufacturing processes [30]; (iv) Digital twin: It integrates data, simulation, and services to create a virtual representation of physical entities, achieving industrial intelligence by utilizing AI, IoT, and ML techniques [27].
@@
While existing methodologies have played an important role in advancing industrial AI, their implementation can be further enhanced by the proposed industrial AI foundation framework. Rather than redefining the core methodologies, the knowledge, data, and model modules offer comprehensive and systematic views that support both existing and emerging industrial AI approaches. By promoting more structured development processes, they ensure that the design, implementation, and deployment of industrial AI solutions are more effective, reliable, and scalable. A concrete example is using ILKM to enable intelligent question-answering (QA) systems for maintenance personnel, who frequently need to query complex maintenance manuals and troubleshooting guides during equipment servicing. The knowledge module enables the structured extraction of key procedures, fault descriptions, and parameter settings from maintenance manuals and technical documentation. This content is organized into knowledge graphs and indexed repositories that can be efficiently queried by LLMs. The data module ensures that domain-specific datasets — including historical maintenance logs and annotated QA pairs — are properly documented, versioned, and linked to the knowledge base. Standardized metadata (e.g., machine type, operating conditions, and fault categories) allows the LLM to retrieve relevant answers aligned with specific equipment and scenarios. The model module guides the fine-tuning of LLMs on domain-specific QA datasets. It also defines best practices for model evaluation and optimization strategies for efficient inference. As a result, maintenance engineers can ask context-aware questions (e.g., ”What should I check if vibration levels exceed the threshold in a [specific] machine?”), and the ILKM-powered QA system provides answers grounded in both structured knowledge and validated historical data.
Figure 3: Case study of applying the industrial artificial intelligence (AI) foundation framework for AI-driven rotating machinery fault diagnosis. Abbreviations: CNN: Convolutional neural network; LLM: Large language model; LSTM: Long short-term memory; ML: Machine learning; RQ: Research question.
@@ 5 Case study
In this section, we demonstrate the application of the proposed industrial AI foundation framework through a case study on intelligent fault diagnosis for rotating machinery. We utilized a publicly available gearbox vibration dataset provided by the Prognostics and Health Management Society [51]. The dataset consists of 2016 labeled samples collected under 78 different operating conditions, covering seven distinct fault classes, including various gear damage levels (class 1, 2, 3, 4, 6, and 8) and healthy state (class 0). Each sample comprises time-domain vibration signals recorded over 3 - 12 seconds at a sampling frequency of 20,480 Hz. The problem addressed in this study is evaluating the multi-class classification performance of the model using this dataset under various operating conditions. Figure 3 illustrates the workflow of applying the industrial AI foundation framework, guiding knowledge extraction, data preparation, model development, and evaluation in a structured and systematic manner.
@@
The process began with the knowledge module. Under its guidance, domain knowledge was extracted from existing publications related to the dataset by leveraging the latest LLMs, GPT-4o, via OpenAI’s application programming interface. Researchers interacted with the LLM using targeted research questions such as: “Can you provide the background information about this dataset?”; “In [target] paper, how did the authors perform data processing?”; and “How can one develop specific ML models such as 1D-CNN, LSTM, or Transformer for this task?” GPT-4o provided structured responses, summaries, and code generation examples, helping researchers quickly understand the dataset’s background, signal characteristics, data preparation requirements, feature extraction strategies, and model development practices.
@@
The data module was applied to ensure systematic preprocessing. A structured pipeline was designed under the guidance of domain knowledge from prior studies. The process began with data cleaning to improve data quality, followed by data segmentation to increase the number of usable samples for model training. Next, feature engineering was performed using the fast Fourier transform to convert raw time-series signals into frequency domain features. The dataset was then split into training, validation, and test sets in a 3:1:1 ratio to ensure robust model evaluation. Throughout this stage, researchers continued to interact with LLMs to validate the soundness of their preprocessing strategies or to quickly obtain code examples for implementing new ideas.
@@
After ensuring that the data were AI-ready, in the model module, researchers developed and evaluated eight different AI models based on previous methodologies and their expertise. These included: (i) tree-based model (decision tree and random forest); (ii) CNN-based model (naive 1D-CNN, residual 1D-CNN); (iii) long short-term memory (LSTM)-based model (naive LSTM, bi-LSTM, hybrid-LSTM); and (iv) transformer-based model (vanilla transformer). Deep learning models were selected for their ability to automatically extract complex patterns from large-scale, frequency-domain features and for their proven robustness in handling variations in operating conditions without the need for extensive manual feature engineering. Hyperparameter tuning and performance evaluation were performed following the experience indicated in the previous research and researchers’ development experience.
@@
The classification accuracy and confusion matrix are presented in Figure 3, demonstrating high performance across various architectures, with the transformer model achieving the highest accuracy of 99.47%. These results highlight the potential of advanced deep learning architectures in industrial fault diagnosis. Compared with previous studies on similar gearbox fault classification tasks, the proposed approach demonstrates clear improvement. For instance, Su and Lee [51] developed a residual CNN that achieved 96.99% accuracy, Vaerenberg et al. [57] used power spectral density preprocessing, log normalization, and a 3-layer CNN to reach 96.9% accuracy, and Gauriat et al. [17] proposed multi-class neural additive models with 92.03% accuracy. The higher performance achieved in this study is attributed to not only the model design but also the systematic application of the industrial AI foundation framework. The structured guidance from the knowledge, data, and model modules ensured consistent data preprocessing, appropriate model selection, and effective hyperparameter tuning, ultimately enhancing reliability, scalability, and real-world applicability.
@@ 6 Future direction
To further strengthen the industrial AI foundation framework, several key directions require further exploration. One critical aspect is talent development. Incorporating 4P-based learning (principle, practice, problem-solving, and professional) and interdisciplinary training in AI/ML, engineering, and industrial applications could benefit the next generation of industrial AI practitioners. Another promising direction is data foundry, which aims to establish a standardized framework for industrial dataset collection, annotation, benchmarking, and management. A well-structured data foundry would enhance collaborative AI research, reproducibility, and cross-industry data sharing while also facilitating the hosting of the Industrial AI Data Challenge Competitions to promote innovation and benchmark AI model performance on industrial datasets. Meanwhile, the foundation framework can be extended toward discrete event dynamic systems and hybrid control systems. This would involve adapting the knowledge, data, and model modules to better handle event-based transitions, symbolic representations, and hierarchical system logic. Furthermore, future research should explore the development of an LLM-assisted intelligent knowledge management system to better make use of historical case studies, domain expertise, and best practices. Such a system could autonomously acquire, structure, and retrieve relevant information, providing researchers and engineers with contextualized and actionable insights to improve AI model development and decision-making in industrial applications.
@@ 7 Conclusion
This paper proposes a unified industrial AI foundation framework, structured into three core modules – the knowledge module, data module, and model module – which collectively support and enhance the industrial AI methodology platform. The role of industrial AI is explored in detail, and a case study on intelligent rotating machinery diagnosis demonstrates the framework’s potential. Furthermore, several future directions are discussed. Overall, our unified industrial AI foundation framework provides a systematic approach to developing, validating, and deploying industrial AI solutions for future industries. Despite these contributions, we acknowledge that acquiring and labeling industrial data remains a significant challenge. Furthermore, while the framework leverages LLMs for knowledge extraction, human guidance is still required, and computational costs may limit its deployment in certain industrial settings. Moving forward, we encourage further research and collaboration between academia, industry, and AI practitioners to refine and expand this framework, fostering the next generation of intelligent, knowledge-data-driven industrial AI ecosystems.